# SuperNeRF-GAN: A Universal 3D-Consistent Super-Resolution Framework for Efficient and Enhanced 3D-Aware Image Synthesis

### v1

@@ III-E2 Loss Function
TABLE II: Effectiveness our proposed SuperNeRF-GAN. Note that all KID scores in this paper are multiplied by 1000.
@@
IV Implementation Details
IV-1 Training Setup
IV-2 SH-HD for Portrait and Cat Face Image Synthesis
V Experiments
V-1 Datasets
V-2 Baselines
For portrait and cat face image synthesis, our model leverages pre-trained EG3D [29] models. We compare our results with GRAM-HD [10], StyleNeRF [27], and StyleSDF [26], all recognized for their high-resolution image synthesis capabilities. For full-body image synthesis, our model builds on the pre-trained SH-HD model. We compare it with EVA3D, VeRi3D, and GSM, which are noted for their 3D-consistent image synthesis.
V-3 Evaluation Metrics
@@
V-A Comparisons
We evaluate 3D-aware image synthesis methods across various criteria, with the results presented in Table I. Note that only our method achieves 3D-consistent HR image synthesis with high universality. Specifically, GRAM-HD is unable to synthesize full-body images due to inherent limitations in its manifolds, and SH-HD struggles with generating high-quality portraits because of boundary issues, which restrict their universality. As our models are deployed on the pre-trained 3D generators of EG3D and SH-HD, we compare our approach with these methods, as illustrated in Table II and LABEL:tab:sh-hd. Compared to EG3D, our method shows stable improvements in 3D-consistency, albeit with a slight compromise in image quality. We argue that the 2D super resolution used in EG3D significantly improve the image quality, whereas SuperNeRF-GAN leverages Depth-Guided Rendering to ensure 3D-consistency. As for SH-HD, which also ensures 3D-consistency, our method outperforms it in both efficiency and image quality. Notably, SH-HD performances poorly on the FFHQ and AFHQ datasets due to its naive depth aggregation technique. Detailed comparisons with other methods are provided below or in the supplementary materials.
V-A1 3D-Consistency
@@ V-A Comparisons
Figure 6: Qualitative comparison of geometry quality. EG3D and our method generate results with more detailed geometry than other methods.
@@ V-A2 Image Quality
In terms of image quality, our method shows significant improvements, as shown in Table V. Notably, although GRAM-HD employs generative radiance manifolds to achieve 3D-consistent image synthesis, it compromises on image quality. In contrast, our method achieves superior performance in both image quality and 3D-consistency. Additionally, we evaluated the image quality of 3D-consistent methods on full-body image synthesis task, with results presented in Table VI. Our method not only achieves the highest image quality but also supports high-resolution image synthesis, thanks to our efficient 3D-consistent super-resolution module.
@@ V-A3 Efficiency
V-A4 Geometry Quality
We conduct a qualitative comparison of geometry quality, as illustrated in Fig. 6. Due to limitations in radiance manifolds, GRAM-HD struggles to produce high-quality geometry results. StyleSDF generates NeRF representations at a lower resolution, which constrains its ability to produce detailed geometry. In contrast, our method integrates seamlessly with existing NeRF-based approaches, resulting in superior geometry quality.
Figure 7: Embedding of images at a resolution of $1024^{2}$ using GAN inversion technique.
@@ V-A4 Geometry Quality
TABLE VIII: Ablation study on interpolation methods. Our proposed Normal-Guided Super-Resolution achieve improvements on 3D-consistency. The results are under $1024^{2}$ resolution to better show the improvements.
@@
V-B Ablation Study
V-B1 Depth Aggregation
The results in Fig. 3 confirm the effectiveness of depth aggregation. We also perform a quantitative ablation study on this technique, presented in Table VII. Specifically, by applying additional dilation and erosion operations, we generate a 5-channel aggregated depth map. The results indicate that doubling the depth aggregation dose not further improve performance, as it introduces depth inaccuracies. Although our Normal-Guided Depth Super-Resolution mitigates this issue, the added complexity does not yield additional benefits.
@@ V-B1 Depth Aggregation
V-B2 Normal-Guided Depth Super-Resolution
To assess the effectiveness of our proposed Normal-Guided Depth Super-Resolution module, we conduct a quantitative comparison with bilinear interpolation. The results in Table VIII show that our method achieves notable improvements in 3D-consistency. This improvement is due to our module’s ability to accurately super-resolve the depth map with the guidance of the normal map, as demonstrated in Fig. 4.
@@ V-C Applications
V-C1 GAN Inversion
Our method can be integrated with existing 3D generative models to embed target images into the latent space of our super-resolved model. This embedding is achieved using GAN inversion techniques, which enable the recovery of high-fidelity images at a resolution of $1024^{2}$, as shown in Fig. 7. Additionally, we explore further applications, such as 3D-consistent image synthesis and style interpolation, which are detailed in the Supplementary Materials.
V-C2 3D-Consistent Image Synthesis
V-C3 High-Resolution Image Synthesis
Fig. 9 showcases high-resolution ($1024\times 1024$) portrait images synthesized using our method. To balance image quality and diversity, we use a truncation technique where the latent code $w$ is defined as a weighted average of two components: $w=0.5\times w_{\mathrm{averaged}}+0.5\times w_{\mathrm{random}}$. Here, $w_{\mathrm{averaged}}$ is the mean latent code across the dataset, ensuring stable and high-quality synthesis, while $w_{\mathrm{random}}$ introduces variability to enhance diversity.
VI Conclusion

### v2

@@ III-E2 Loss Function
IV Implementation Details
IV-1 Training Setup
IV-2 SH-HD for Portrait and Cat Face Image Synthesis
V Experiments
V-1 Datasets
V-2 Baselines
For portrait and cat face image synthesis, our model leverages pre-trained EG3D [29] models. We compare our results with GRAM-HD [10], StyleNeRF [27], and StyleSDF [26], all recognized for their high-resolution image synthesis capabilities. However, StyleNeRF and StyleSDF rely on 2D super-resolution, compromising the 3D consistency of the synthesized images. In contrast, GRAM-HD’s manifold representation limits image quality, geometry accuracy, and universality. For full-body image synthesis, our model builds on the pre-trained SH-HD model. We compare it with EVA3D, VeRi3D, and GSM, which are noted for their 3D-consistent image synthesis. Nevertheless, their reliance on vertex-based, MLP-based, or Gaussian shell-based representations imposes limitations on the quality of the generated images.
V-3 Evaluation Metrics
TABLE II: Effectiveness of SuperNeRF-GAN for portrait image synthesis. Note that all KID scores in this paper are multiplied by 1000.
@@ V-3 Evaluation Metrics
V-A Comparisons
We evaluate 3D-aware image synthesis methods across various criteria, with the results presented in Table I. Note that only our method achieves 3D-consistent HR image synthesis with high universality. Specifically, GRAM-HD is unable to synthesize full-body images due to inherent limitations in its manifolds, and SH-HD struggles with generating high-quality portraits because of boundary issues, which restrict their universality.
As our models are deployed on the pre-trained 3D generators of EG3D and SH-HD, we compare our approach with these methods, as illustrated in Table II and LABEL:tab:sh-hd. Compared to EG3D, our method shows stable improvements in 3D-consistency, albeit with a slight compromise in image quality. We argue that the 2D super resolution used in EG3D significantly improve the image quality, whereas SuperNeRF-GAN leverages Depth-Guided Rendering to ensure 3D-consistency. It is worth noting that the quantitative improvements in 3D-consistency may appear subtle, as they are constrained by the limitations of the reconstruction method NeuS. We encourage readers to refer to the Supplementary Video for more qualitative comparisons, where the advantages of our method are clearly evident. As for SH-HD, which also ensures 3D-consistency, our method outperforms it in both efficiency and image quality. Notably, SH-HD performances poorly on the FFHQ and AFHQ datasets due to its naive depth aggregation technique.
V-A1 3D-Consistency
@@ V-A Comparisons
Figure 6: Qualitative comparison of geometry quality. EG3D and our method generate results with more detailed geometry than other methods. The results for GRAM-HD, StyleNeRF, and StyleSDF are directly taken from their respective papers.
@@ V-A2 Image Quality
In terms of image quality, our method shows significant improvements, as shown in Table V. Notably, although GRAM-HD employs generative radiance manifolds to achieve 3D-consistent image synthesis, it compromises on image quality compared to super-resolution-based methods such as StyleNeRF and StyleSDF. This trade-off highlights that achieving strong 3D-consistency imposes stronger constraints on generators, often leading to a degradation in image quality. Consequently, we argue that the slight compromise in image quality in our method, as compared to EG3D (Table II), is acceptable. On the other hand, while our method prioritizes 3D-consistency, it still achieves better image quality than StyleNeRF and StyleSDF, and is comparable to EG3D. Furthermore, we evaluated the image quality of 3D-consistent methods on full-body image synthesis task, with results presented in Table VI. Our method not only achieves the highest image quality but also supports high-resolution image synthesis with low memory consumption, thanks to our efficient 3D-consistent super-resolution module.
@@ V-A3 Efficiency
V-A4 Universality
As noted in the GRAM-HD paper, this method cannot synthesize images from wide viewpoints due to its reliance on the near-plane manifold representation for efficient rendering. In contrast, our method leverages depth-guided rendering, which is compatible with any NeRF representation and enables the synthesis of images from wide viewpoints, such as full-body images. Another 3D-consistent method, SH-HD, struggles with unsmooth depth values, as shown in Table LABEL:tab:sh-hd, due to its naive neighbor-aware depth aggregation technique. This issue is further evidenced in Fig.3. Consequently, our method stands as the only 3D-consistent image synthesis approach with robust universality, as demonstrated in TableI.
V-A5 Geometry Quality
We conduct a qualitative comparison of geometry quality, as shown in Fig. 6. GRAM-HD employs radiance manifolds to achieve efficient rendering, with each manifold being nearly planar. As a result, the generated geometries are essentially the combination of several ”planes,” which limits the overall quality. StyleSDF and StyleNeRF generate NeRF representations at relatively low resolutions (e.g., 32), which further constrains their ability to capture fine geometric details. In contrast, our method produces high-resolution NeRF representations, achieving the same geometry quality as EG3D.
@@ V-A5 Geometry Quality
V-B Ablation Study
V-B1 Depth Aggregation
The results in Fig.3 confirm the effectiveness of our depth aggregation technique, as discussed in SectionIII-D. Specifically, without depth aggregation, noticeable holes appear in the generated images due to the inaccuracies caused by straightforward bilinear interpolation. In contrast, the naive neighbor-aware depth aggregation proposed in SH-HD suffers from artifacts at depth discontinuities. This approach only partially mitigates the boundary issue without fundamentally resolving it. This claim is further supported quantitatively in Table LABEL:tab:sh-hd, where our method demonstrates greater universality compared to SH-HD, primarily due to the differences in depth aggregation techniques.
To further investigate depth aggregation, we conduct a quantitative ablation study, as presented in Table VII. In this study, we apply additional dilation and erosion operations to produce a 5-channel aggregated depth map. The results show that doubling the depth aggregation does not improve performance and instead introduces depth inaccuracies. While our Normal-Guided Depth Super-Resolution mitigates these inaccuracies, the additional depth values fail to yield benefits.
TABLE VIII: Ablation study on interpolation methods. Our proposed Normal-Guided Super-Resolution achieve improvements on 3D-consistency. The models are trained on the FFHQ1024 dataset, as high resolution better highlights these improvements.
@@ V-B1 Depth Aggregation
V-B2 Normal-Guided Depth Super-Resolution
To evaluate the effectiveness of our proposed Normal-Guided Depth Super-Resolution module, we perform a quantitative comparison against bilinear interpolation. As shown in Table VIII, our method achieves notable improvements in 3D-consistency. These improvements stem from the module’s ability to accurately super-resolve depth maps with guidance from the normal map. As illustrated in Fig.4, compared to naive bilinear interpolation, our normal-guided approach leverages the normal map to identify and handle depth discontinuities at boundaries more effectively. This capability ensures smoother and more accurate depth transitions. For a detailed explanation, please refer to SectionIII-D2.
Figure 7: Interpolation in the latent space produces smooth transitions, with each intermediate result maintaining 3D-consistency.
@@ V-C Applications
V-C1 3D-Consistent Image Synthesis
V-C2 Interpolation in Latent Space
Fig. 7 demonstrates the results of interpolation in latent space. Specifically, given two random latent codes, we interpolate between them and feed the interpolated codes into the generator. This process produces smooth transitions in the synthesized images. Notably, each interpolated result exhibits 3D consistency, demonstrating that images synthesized across the entire latent space can maintain 3D consistency, thanks to our direct rendering strategy.
V-C3 High-Resolution Image Synthesis
Fig. 9 showcases high-resolution portraits synthesized by our method. To balance image quality and diversity, we use a truncation technique where the latent code $w$ is defined as a weighted average of two components: $w=0.5\times w_{\mathrm{averaged}}+0.5\times w_{\mathrm{random}}$. Here, $w_{\mathrm{averaged}}$ is the averaged latent code, ensuring high-quality synthesis, while $w_{\mathrm{random}}$ introduces diversity. For each generated entity, we provide images from different viewpoints, demonstrating the 3D-consistency of high-resolution image synthesis—a challenge for previous methods.
@@ V-C3 High-Resolution Image Synthesis
VI limiation
Our method achieves stronger 3D consistency than previous approaches by directly rendering high-resolution images. However, this also results in a slight degradation in image quality compared to methods based on 2D super-resolution. Exploring approaches that balance both strong 3D consistency and high image quality would be valuable for broader industrial applications. Additionally, while our method enables efficient synthesis of high-resolution images through depth-guided rendering, it still requires dense sampling during the low-resolution stage, which can hinder real-time rendering. Replacing NeRF with 3DGS presents a promising alternative, though this area has been explored relatively less.
VII Conclusion